OpenStudy (anonymous):

36-(8+4) x2-3 a. 6 b. 9 c. 33 d. 45

OpenStudy (anonymous):

36-12x^2-3? if that is set equal to zero. 0=36-12x^2-3 -33=-12x^2 -33/-12= x^2 \[\sqrt{-33/-12}\] = 1.685 = x

OpenStudy (anonymous):

look at the choices

OpenStudy (anonymous):

Partygirl, is it x2-3 or *2-3?

OpenStudy (anonymous):

You have to be careful with how you state the question, If its * instead of x, then your answer is b, 9

OpenStudy (precal):

use order of operations

OpenStudy (anonymous):

how you get 9

OpenStudy (precal):

36-(8+4) x2-3 36-(12)x2-3 36-24-3 12-3 9

OpenStudy (precal):

Order of operation do ( ) first multiply subtract left to right

Directrix (directrix):

^^^ Precal's work above is correct - I concur.

OpenStudy (anonymous):

explain me

OpenStudy (precal):



OpenStudy (precal):

P is for parenthesis (always do what is inside first) E is for exponents MD is multiplication and Division do it left to right (whoever is first) example 8/4X3 = 2X3= 6 AS is addition and subtraction do it left to right (whoever is first) example 1-3+3=-2+3=1

Directrix (directrix):

@partyrainbow276 --> 36 - (8+4) *2 - 3 = Do the ( ) first. 36 - (12) * 2 - 3 = Do the multiplication next --> (12) * 2 36 - 24 - 3 = Do subtraction next 12 - 3 = 9

OpenStudy (anonymous):

I did it but i got 5.

OpenStudy (anonymous):

Can you show your procedure? Maybe we can spot the problem

OpenStudy (anonymous):

She is explaining this wrong as pellet.

OpenStudy (anonymous):

i did 8+4=12 then 12x2= 24 then 24-36= 8 then -3 and got 5

OpenStudy (anonymous):

and now im doing it again by myself

OpenStudy (anonymous):

Thats where you are getting it wrong

OpenStudy (anonymous):

you see 36-24 = 12 not 8 then, you substract by 3

OpenStudy (anonymous):

=9

Directrix (directrix):

@partyrainbow276 i did 8+4=12 OK then 12x2= 24 OK then 24-36= 8 NOT OK --> Look at the problem: It is 36 - 24 which is 12. then 12 -3 and you will get 9

OpenStudy (anonymous):

I GOT IT!!!! WOOO Thx open2study and my friend directrix

OpenStudy (anonymous):

thx for taking out the time of ur day helping me

Directrix (directrix):

You might want to practice on another one.

OpenStudy (anonymous):

Yea. 0-0

OpenStudy (anonymous):

Anytime:)

OpenStudy (anonymous):

check this one -15 bangor -18 bichmark 9 los angeles 15 miami hawain 19 Is this in right order? of integers

Directrix (directrix):

-18 is smaller than -15. On a number line, the smaller number lives to the left of the larger. -18 is to the left of -15 so -18 is smaller. 18 bichmark- 15 bangor - 9 los angeles 15 miami hawain 19 is how it goes.

OpenStudy (anonymous):

Another medal for directrix:)

Directrix (directrix):

@Open2study --> not in this thread